OpenStudy (anonymous):

Show that \[a_n=\sin(n)n^{-1/2}\] and \[b_n=\sin(\frac{ n \pi }{ 2 }+\frac{ \pi }{ 3 })n^{-1/2}\] converges or diverges

OpenStudy (anonymous):

I assume that it is convergent. Now I have:\[\sin(n)n^{-1/2}=\frac{ \sin(n) }{ \sqrt(n) }\] And then I want to use the definition for convergence, so I let \[\epsilon >0\] and now I want to find a \[N \in \mathbb{N}:|a_n-a|\le \epsilon \]. From here I am having rouble proving or dis-proving this.

OpenStudy (anonymous):

Does it make sens to split \[a_n\] into 2: \[sin(n), n^{-1/2}\] And then I could do: \[n^{-1/2}=1/sqrt(n)<1/n<epsilon\] Now here we know 1/n converges towards 0, hereby we see that \[n^{-1/2}\] also converges towards 0. And since \[sin(n)\] is limited: \[-1<=sin(n)<=1\] The whole sequence is convergent towards 0? - using the multiplication rule for sequences.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ n^{-1/2}\sin(n) }\) \(\large\color{black}{ \displaystyle \sin(n) =\displaystyle \sum_{ i=1 }^{ \infty }\frac{(-1)^kn^{1+2k}}{(1+2k)!} }\) \(\large\color{black}{ \displaystyle n^{-1/2}\sin(n) =\displaystyle \sum_{ i=1 }^{ \infty }\frac{(-1)^kn^{1+2k~~-1/2 }}{(1+2k)!} }\) \(\large\color{black}{ \displaystyle n^{-1/2}\sin(n) =\displaystyle \sum_{ i=1 }^{ \infty }\frac{(-1)^kn^{2k+1/2 }}{(1+2k)!} }\) \(\large\color{black}{ \displaystyle \int n^{-1/2}\sin(n) {\rm ~d}n=\displaystyle \sum_{ i=1 }^{ \infty }\frac{(-1)^kn^{2k+3/2 }}{(2k+3/2)(1+2k)!} }\)

OpenStudy (solomonzelman):

So the first series converges byu the integra test

OpenStudy (solomonzelman):

Correction: (grammar) integral test**

OpenStudy (anonymous):

I haven't learned about the integral test yet. I know the definition for convergence, and cauchy sequence.

OpenStudy (math&ing001):

Use the Riemann Series: \[a _{n}=\frac{ 1 }{ n ^{q} }\] The series is: Divergent when q<=1 Convergent when q>1

OpenStudy (solomonzelman):

Wel, the integral test is that you integrate A_n. If the integral converges, then series also converges. If integral diverges, then series also diverges. (You can take integral with limits ∞ and 1, or ∞ and 100, doesn't matter)

OpenStudy (solomonzelman):

Loser66, I checked the sries A_n does converge. http://www.wolframalpha.com/input/?i=summation+from+n%3D1+to+n%3D%E2%88%9E+of+sin%28n%29%2F%E2%88%9An

OpenStudy (math&ing001):

Ignore that, not gonna work here. Your series has to be inferior to a converging series, but here 1/sqrt(n) is divergent.

OpenStudy (anonymous):

Doesnt in converge towards 0? Since \[1/sqrt(n)\] converges towards 0, and then if you multiply it with a limited sequence: \[-1<=sin(n)<=1\] You would get 0

OpenStudy (math&ing001):

the series 1/sqrt(n) is not convergent

OpenStudy (solomonzelman):

it meets alternating series test conceptualy....

OpenStudy (solomonzelman):

Terms decrease, there is a non-integer alternation, if you want.... Try partial sums to confirm

OpenStudy (solomonzelman):

But without the sin(n), it would indeed diverge, yes.

OpenStudy (solomonzelman):

because sin(n)/n is not 1 as n->∞

OpenStudy (solomonzelman):

it is 0

OpenStudy (solomonzelman):

yes:)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{n=1}^{\infty}\sin\left(\frac{ n \pi }{ 2 }+\frac{ \pi }{ 3 }\right) n^{-1/2} }\) \(\large\color{black}{ \displaystyle \int_{n=1}^{n=\infty} n^{-1/2}\sin\left(\frac{ n \pi }{ 2 }+\frac{ \pi }{ 3 }\right) {\rm ~d}n=}\) \(\large\color{blue}{ \displaystyle u=\left(\frac{ n \pi }{ 2 }+\frac{ \pi }{ 3 }\right) }\) \(\large\color{blue}{ \displaystyle du=\frac{ \pi }{ 2 } dn\quad \Longrightarrow\frac{ 2 } { \pi } du=dn}\) \(\large\color{blue}{ \displaystyle n=1\quad \Longrightarrow u=\left(\frac{ \color{red}{1\cdot } \pi }{ 2 }+\frac{ \pi }{ 3 }\right)=\frac{5\pi}{6}}\) \(\large\color{blue}{ \displaystyle n=\infty \quad \Longrightarrow u=\left(\frac{ \color{red}{\infty\cdot } \pi }{ 2 }+\frac{ \pi }{ 3 }\right)=\infty}\) \(\large\color{black}{ \displaystyle \frac{ 2 } { \pi } \int_{u=\frac{5\pi}{6}}^{u=\infty} n^{-1/2}\sin\left(u\right) {\rm ~d}u=}\)

OpenStudy (solomonzelman):

By the same tocken, the second series goes by the integral test. Converges :)

OpenStudy (solomonzelman):

(because the integral would converge 9since it is even smaller than the integral by the A_n)

OpenStudy (solomonzelman):

Just in case, here is a link to integral test. One possible resource for it. http://tutorial.math.lamar.edu/Classes/CalcII/IntegralTest.aspx

OpenStudy (anonymous):

Thanks, ill get a look at it

OpenStudy (solomonzelman):

Sure enjoy ...

OpenStudy (solomonzelman):

One more correction, in my last line of the first reply, the "i=1" index under the sigma is supposed to be "k=1".

jagr2713 (jagr2713):

Great job guys :D Thanks for helping